OpenStudy (anonymous):

WILL GIVE BEST RESPONSE Use mathematical induction to prove the statement is true for all positive integers n. 10 + 30 + 60 + ... + 10n = 5n(n + 1)

OpenStudy (amistre64):

step 1, is it true for n=1?

OpenStudy (amistre64):

if so, assume its true for n=k and then add k+1 to the mix and see if it simplifies to an n=k+1 model

OpenStudy (anonymous):

can you show me the process please?

OpenStudy (amistre64):

step 1: does 10(1) = 5(1)(1+1) ?

OpenStudy (anonymous):

yes it does

OpenStudy (amistre64):

then we have a basis step, so lets assume its true for any n=k and write 10 + 30+ ... + 10k = 5k(k + 1) agreed?

OpenStudy (anonymous):

yes I agree

OpenStudy (amistre64):

now, the next step lets us know that if we add (k+1), then by induction its proved

OpenStudy (amistre64):

lets add (k+1), but i might have said that wrong, lets add one more element to the left side; and add the same amount to the right side 10 + 30+ ... + 10k + 10(k+1) = 5k(k + 1) + 10(k+1)

OpenStudy (anonymous):

okay I understand that part

OpenStudy (amistre64):

im trying to make sure i understand it too :) now, if we can get the right side to equal: 5(k+1)(k+1+1) its proved

OpenStudy (amistre64):

\[5k(k + 1) + 10(k+1)\]factor out a (k+1) \[(k+1)(5k+10)\]factor our a 5 \[5(k+1)(k+2)\]and rewrite k+2 as k=1+1 \[5(k+1)((k+1)+1)\] tada!!

OpenStudy (amistre64):

now, we know that it works for k=1 and therefore we can be assured that it works for k=2, k=3, k=4, ...

OpenStudy (anonymous):

Thank you so much! You made it seem so much easier to me, I really appreciate it!

OpenStudy (amistre64):

youre welcome, i had to dig deep to remember that one :)